Two vectors 10N and 8N of bearing 285° and N70°E respectively. Find the resultant force using triangle of vectors.
1
Expert's answer
2019-08-19T09:38:26-0400


using cosine rule - "F^{2}_{R}=8^{2}+10^{2}-2{\cdot}(8{\cdot}10){\cdot}cos{35^{o}}=5.74" using sine rule, we can find the angle of bearing

"\frac{sina}{10}=\frac{sin35^{o}}{5.74} \to 86.4^{o}"